Computing and Assessing Plant Asset Impairment Zeibart Company purchases equipment for $225,000 on July 1, 2016, with an estiZeibart Company purchases equipment for $225,000 on July 1, 2016, with an estimated useful life of 10 years and expected salvage value of $25,000. Straight-line depreciation is used. On July 1, 2020, economic factors cause the market value of the equipment to decline to $90,000. On this date, Zeibart examines the equipment for impairment and estimates $125,000 in future cash inflows related to use of this equipment.

a. Is the equipment impaired at July 1, 2020?
AnswerYesNo

b. If the equipment is impaired on July 1, 2020, compute the impairment loss and prepare a journal entry to record the loss. If equipment is not impaired, select "N/A" as your Description answers and leave the Debit and Credit answers blank (zero).

c. What amount of depreciation expense would Zeibart record for the 12 months from July 1, 2020 through June 30, 2021? Prepare a journal entry to record this depreciation expense. Round to the nearest dollar. (Hint: Assume no change in salvage value.)

d. Using the financial statement effects template, show how the entries in parts b and c affect Zeibart Company’s balance sheet and income statement.
Answer #1

a.

Net book value = Cost - (Cost - Salvage value) / Estimated useful life x Period covered

                                = $225,000 - ($225,000 - $25,000) / 10 years x 4 years = $ 145,000

The net book value is greater than the recoverable amount of the equipment. Hence, the equipment is impaired.

b.

The recoverable amount is higher of fair market value or future cash flows

Recoverable amount = $90,000 Vs $125,000 = $125,000

Impairment loss = Net book value - Recoverable amount = $145,000 - $125,000 = $20,000

The new value of the equipment is $125,000 ($225,000 - $100,000).

The required journal entry would be

Date

Particulars

Debit ($)

Credit ($)

Impairment loss

20,000

Accumulated depreciation

80,000

           Equipment

100,000

(To record impairment)

c.

Depreciation      = New value of the equipment / Remaining estimated useful life

=$125,000 / (10 years - 4 years) = $20,833

Date

Particulars

Debit ($)

Credit ($)

Depreciation expense

20,833

             Accumulated depreciation

20,833

(To record the depreciation)

d.

Balance sheet

Transaction

Cash asset

+ noncash
asset

-Contra assets

= liabilities

+ contri
capital

+ earned
capital

Impairment charges

(100,000)

-(80,000)

-(20,000)

(100,000)

Depreciation expense

(20,833)

-(20833)

(20,833)

Note that these impairment loss and depreciation will reduce the profit and loss and ultimately earned capital.

Income statement

Transaction

Revenue

-Income

= net income

Impairment charges

-20,000

-80,000

=(100,000)

Depreciation expense

-20,833

=(20,833)